# Formatting Instructions for ICLR 2024 Conference Submissions

### v3

@@ 1 Introduction
Numerous methods have been developed to estimate fitness functions from experimental data, including classical machine learning techniques [42], deep learning approaches [19], and semi-supervised methods [21]. Additionally, there are many methods that incorporate biological assumptions into the modeling process, such as parameterized biophysical models [28], non-parametric techniques [43, 44], and methods for spectral regularization of neural networks [1]. These latter approaches largely focus on accurately modeling the influence of “epistasis" on fitness functions, which refers to statistical or physical interactions between genetic elements, typically either amino-acids in a protein sequence or genes in a genome.
“Local" epistasis refers to interactions between a limited number of specific positions in a sequence, and is often modeled using interaction terms in a linear model of a fitness function [29]. “Global" epistasis, on the other hand, refers to the presence of nonlinear relationships that affect the fitness of sequences in a nonspecific manner. A model of global epistasis typically assumes a simple latent fitness function is transformed by a monotonically increasing nonlinearity to produce observed fitness data [35, 30, 38, 33, 36]. Typically, these models assume a particular parametric form of the latent fitness function and nonlinearity, and fit the parameters of both simultaneously. It is most common to assume that the underlying fitness function includes only additive (non-interacting) effects [30], though pairwise interaction effects have been added in some models [38].
@@ 2.1 Fitness functions and the Graph Fourier transform
where ${\bm{f}}$ is a length $q^{L}$ vector containing the fitness values of every sequence in $\mathcal{S}$, ${\bm{\Phi}}$ is a $q^{L}\times q^{L}$ orthogonal matrix representing the Graph Fourier basis, and $\bm{\beta}$ is a length $q^{L}$ vector containing the weights corresponding to all possible epistatic interactions. We refer to ${\bm{f}}$ and $\bm{\beta}$ as representing the fitness function in the fitness domain and the epistatic domain, respectively. Note that we may apply the inverse transformation of Eq.  1 to any complete observed fitness function, ${\bm{y}}$ to calculate the epistatic representation of the observed data, $\bm{\beta}_{{\bm{y}}}={\bm{\Phi}}^{T}{\bm{y}}$. Similarly, if $\hat{{\bm{f}}}$ contains the predictions of a fitness model for every sequence in a sequence space, then $\hat{\bm{\beta}}={\bm{\Phi}}^{T}\hat{{\bm{f}}}$ is the epistatic representation of the model.
A fitness function is considered sparse, or concentrated, in the epistatic domain when $\bm{\beta}$ contains a relatively small number of elements with large magnitudes, and many elements equal to zero or with small magnitudes. In what follows, we may refer to a fitness function that is sparse in the epistatic domain as simply being a “sparse fitness function". A number of experimentally-determined fitness functions have been observed to be sparse in the epistatic domain [31, 18, 7]. Crucially, the sparsity of a fitness function in the epistatic domain determines how many measurements are required to estimate the fitness function using Compressed Sensing techniques that minimize a MSE loss function [7]. Herein we consider the effect that global epistasis has on a sparse fitness function. In particular, we argue that global epistasis results in observed fitness functions that are dense in the epistatic domain, and thus require a large amount of data to accurately estimate by minimizing a MSE loss function. However, in these cases, there may be a sparse ranking function that can be efficiently extracted by minimizing a contrastive loss function.
@@ 2.2 Global epistasis models
Global epistasis models of the form of Eq.  2 have proved effective at capturing the variation observed in empirical fitness data [26, 36, 30, 32, 33, 38], suggesting that global epistasis is a common feature of natural fitness functions. Further, it has been shown that global epistasis results from first-principles physical considerations that are common in many biological systems. In particular, Husain & Murugan [22] show that global epistasis arises when the physical dynamics of a system is dominated by slow, collective modes of motion, which is often the case for protein dynamics. Aside from intrinsic/endogenous sources, the process of measuring fitness can also introduce nonlinear effects that are dependent on the experiment and not on the underlying fitness function. For example, fitness data is often left-censored, as many sequence have fitness that falls below the detection threshold of an assay. Finally, global diminishing-returns epistatic patterns have been observed widely in both single and multi-gene settings where the interactions are among genes rather than within a gene [26, 33, 4].
Together, these results indicate that global epistasis is an effect that can be expected in empirically-observed fitness functions. In what follows, we argue that global epistasis is detrimental to effective modeling of fitness functions using standard techniques. In particular, global epistasis manifests itself by producing observed data that is dense in the epistatic domain. In other words, when observed fitness data is produced through Eq.  2 then the epistatic representation of this fitness function (calculated through application of Eq.  1), is not sparse. Further we argue that this effect of global epistasis makes it to difficult to model such observed data by minimizing standard MSE loss functions with a fixed amount of data. Further, we argue that fitting fitness models aimed at extracting the latent fitness function from observed data is a more efficient use of observed data that results in improved predictive performance (in the ranking sense).
@@ 2.3 Contrastive losses
where $f_{{\bm{\theta}}}$ is a model with parameters ${\bm{\theta}}$, ${\bm{x}}_{i}$ are model inputs and $y_{i}$ are the corresponding labels of those inputs. This loss compares every pair of data points and encourages the model output $f_{{\bm{\theta}}}({\bm{x}}_{i})$ to be greater than $f_{{\bm{\theta}}}({\bm{x}}_{j})$ whenever $y_{i}>y_{j}$; in other words, it encourages the model outputs to be ranked according to their labels. A number of distinct but similar loss functions have been proposed in the learning-to-rank literature [13] and also for metric learning [14]. An example is the Margin ranking loss [20], which replaces the logistic function in the sum of Eq.  3 with a hinge function.In our experiments, we largely focus on the BT loss of Eq.  3 as we found it typically results in superior predictive performance; however, we also compare to models trained with the Margin loss when presenting results in benchmark tasks.
@@
A key feature of the contrastive loss in Eq.  3 is that it only uses information about the ranking of observed labels, rather than the numerical values of the labels. Thus, the loss is unchanged when the observed values are transformed by a monotonic nonlinearity. We will show that this feature allows this loss to recover a sparse latent fitness function from observed data that has been affected by global epistasis, and enables more data-efficient learning of fitness functions compared to a MSE loss.
@@ 3 Results
Our results are aimed at demonstrating three properties of contrastive losses. First, we show that given complete, noiseless fitness data (i.e. noiseless fitness values associated with every sequence in the sequence space) that has been affected by global epistasis, minimizing the BT loss enables a model to nearly exactly recover the sparse latent fitness function $f$. Next, we consider the case of incomplete data, where the aim is to predict the relative fitness of unobserved sequences. In this regime, we find through simulation that minimizing the BT loss enables models to achieve better predictive performance then minimizing the MSE loss when the observed data has been affected with global epistasis. We argue by way of a fitness-epistasis uncertainty principle that this is due to the fact that nonlinearities tend to produce fitness functions that do not admit a sparse representation in the epistatic domain, and thus require more data to learn with MSE loss. Finally, we demonstrate the practical significance of these insights by showing that minimizing the BT loss results in improved predictive performance over MSE loss in nearly all tasks in the FLIP benchmark [15].
@@ 3.1 Recovery from complete data
We first consider the case of “complete" data, where fitness measurements are available for every sequence in the sequence space. The aim of our task in this case is to recover a sparse latent fitness function when the observed measurements have been transformed by an arbitrary monotonic nonlinearity. In particular, we sample a sparse fitness function $f$ from the NK model [24], a popular model of fitness functions that has been shown to recapitulate the sparsity properties of some empirical fitness functions [7]. The NK model has three parameters: $L$, the length of the sequences, $q$, the size of the alphabet for sequence elements, and $K$, the maximum order of (local) epistatic interactions in the fitness function. Roughly, the model randomly assigns $K-1$ interacting positions to each position in the sequence, resulting in a sparse set of interactions in the epistatic domain. The weights of each of the assigned interactions are then drawn from a independent unit normal distributions.
@@
We then transform the sampled fitness function $f$ with a monotonic nonlinearity $g$ to produce a set of complete observed data, $y_{i}=g(f({\bm{x}}_{i}))$ for all ${\bm{x}}_{i}\in\mathcal{S}$. The goal of the task is then to recover the function $f$ given all $({\bm{x}}_{i},y_{i})$ pairs. In order to do so, we model $f$ using a two layer fully connected neural network and fit the parameters of this model by performing stochastic gradient descent (SGD) on the BT loss, using the Spearman correlation between model predictions and the $y_{i}$ values to determine convergence of the optimization. We then compare the resulting model, $\hat{f}$, to the latent fitness function $f$ in both the fitness and epistatic domains, using the forward and inverse transformation of Eq.  1 to convert between the two domains.
@@
Fig. 1 shows the results of one of these tests. In this case, we used an exponential function to represent the global epistasis nonlinearity. The exponential function exaggerates the effects of global epistasis in the epistatic domain and thus better illustrates the usefulness of contrastive losses, although the nonlinearities in empirical fitness functions tend to have a more sigmoidal shape [30]. Fig. 1b shows that the global epistasis nonlinearity substantially alters the representations of the observed data $y$ in both the fitness and epistatic domains, as compared to the latent fitness function $f$. Nonetheless, Fig. 1c demonstrates that the model fitness function $\hat{f}$ created by minimizing the BT loss is able to nearly perfectly recover the sparse latent fitness function (where recovery is defined as being equivalent up to an affine transformation). This is a somewhat surprising result, as there are many fitness functions that correctly rank the fitness of sequences, and it is not immediately clear why minimizing the BT loss produces this particular sparse latent fitness function. However, this example makes clear that fitting a model by minimizing the BT loss can be an effective strategy for recovering a sparse latent fitness function from observed data that has been affected by global epistasis. Similar results from additional examples of this task using different nonlinearities and latent fitness functions are shown in Appendix B.1. Further, in Appendix B.2 we compare the results of Fig. 1c to a baseline method in which a quantile transformation is applied to the observed data.
@@ 3.2 Fitness-epistasis uncertainty principle
Figure 2: (a) Demonstration of the fitness-epistasis uncertainty principle where a latent fitness function is transformed by $g(f)=\exp(a\cdot f)$ for various settings of $a$. The dashed black line indicates the lower bound on the sum of the entropies of the fitness and epistatic representations of the fitness function (b) Test-set Spearman correlation for models trained with MSE and BT loss on incomplete fitness data transformed by various nonlinearities, compared to the entropy of the fitness function in the epistatic domain. Each point corresponds to a model trained on 256 randomly sampled training points from a $L=10,K=2$ latent fitness function which was then transformed by a nonlinearity. (c) Convergence comparison between models fit with BT and MSE losses to observed data generated by transforming an $L=10,K=2$ latent fitness function by $g(f)=\exp(10\cdot f)$. Each point represents the mean test set correlation over 200 training set replicates.
@@
Fig. 1b provides an example where a monotonic nonlinearity applied to a sparse fitness increases the density of the the fitness function in the epistatic domain. In particular, we see that many “spurious" local epistatic interactions must appear in order to represent the nonlinearity (e.g. interactions of order 3, when we used an NK model with $K=2$). This effect can be observed for many different shapes of nonlinearities [35, 3]. We can understand this effect more generally using uncertainty principles, which roughly show that a function cannot be concentrated on a small number of values in two different representations. In particular, we consider the discrete entropic uncertainty principle proved by Dembo et al. [16]. When applied to the transformation in Eq.  1, this uncertainty principle states:
$H({\bm{f}})+H(\bm{\beta})\geq L\log\left(\dfrac{1}{m^{2}}\right),$ (4)
where $H({\bm{x}})=-\sum_{i}\frac{x^{2}_{i}}{||x||^{2}}\log\frac{x^{2}_{i}}{||x||^{2}}$ is the entropy of the normalized squared magnitudes of a vector and $m=1/\sqrt{q}$ when $q=2$, $m=1/(q-\sqrt{q})$ when $q=3$ and $m=1-1/(q-\sqrt{q})$ otherwise. Low entropy indicates that a vector is highly concentrated on a small set of elements. Thus, the fitness-epistasis uncertainty principle of Eq.  4 shows that fitness functions cannot be highly concentrated in both the fitness and epistatic domains. A sparse fitness function (in the epistatic domain) must therefore be “spread out" (i.e. dense) in the fitness domain, and vice-versa.
The importance of this result for understanding global epistasis is that applying a nonlinearity to a dense vector will often have the effect of concentrating the vector on a smaller number of values. This can most easily be seen for convex nonlinearities like the exponential shown in Fig 1a, but is also true of many other nonlinearities (see Appendix C: for additional examples). If this concentration in the fitness domain is sufficiently extreme, then the epistatic representation of the fitness function, $\bm{\beta}$, must be dense in order to satisfy Eq.  4. Fig 2a demonstrates the uncertainty principle by showing how the entropy in the fitness and epistatic domains decrease and increase, respectively, as more extreme nonlinearities are applied to a sparse latent fitness function.

### v4

# Contrastive losses as generalized models of global epistasis

@@ 1 Introduction
Numerous methods have been developed to estimate fitness functions from experimental data, including classical machine learning techniques [43], deep learning approaches [18], and semi-supervised methods [21]. Additionally, there are many methods that incorporate biological assumptions into the modeling process, such as parameterized biophysical models [29], non-parametric techniques [44, 45], and methods for spectral regularization of neural networks [1]. These latter approaches largely focus on accurately modeling the influence of “epistasis” on fitness functions, which refers to statistical or physical interactions between genetic elements, typically either amino-acids in a protein sequence or genes in a genome.
“Local” epistasis refers to interactions between a limited number of specific positions in a sequence, and is often modeled using interaction terms in a linear model of a fitness function [30]. “Global” epistasis, on the other hand, refers to the presence of nonlinear relationships that affect the fitness of sequences in a nonspecific manner. A model of global epistasis typically assumes a simple latent fitness function is transformed by a monotonically increasing nonlinearity to produce observed fitness data [36, 31, 39, 34, 37]. Typically, these models assume a particular parametric form of the latent fitness function and nonlinearity, and fit the parameters of both simultaneously. It is most common to assume that the underlying fitness function includes only additive (non-interacting) effects [31], though pairwise interaction effects have been added in some models [39].
@@ 2.1 Fitness functions and the Graph Fourier transform
where ${\bm{f}}$ is a length $q^{L}$ vector containing the fitness values of every sequence in $\mathcal{S}$, ${\bm{\Phi}}$ is a $q^{L}\times q^{L}$ orthogonal matrix representing the Graph Fourier basis, and $\bm{\beta}$ is a length $q^{L}$ vector containing the weights corresponding to all possible epistatic interactions. We refer to ${\bm{f}}$ and $\bm{\beta}$ as representing the fitness function in the fitness domain and the epistatic domain, respectively. Note that we may apply the inverse transformation of Eq. 1 to any complete observed fitness function, ${\bm{y}}$ to calculate the epistatic representation of the observed data, $\bm{\beta}_{{\bm{y}}}={\bm{\Phi}}^{T}{\bm{y}}$. Similarly, if $\hat{{\bm{f}}}$ contains the predictions of a fitness model for every sequence in a sequence space, then $\hat{\bm{\beta}}={\bm{\Phi}}^{T}\hat{{\bm{f}}}$ is the epistatic representation of the model.
A fitness function is considered sparse, or concentrated, in the epistatic domain when $\bm{\beta}$ contains a relatively small number of elements with large magnitudes, and many elements equal to zero or with small magnitudes. In what follows, we may refer to a fitness function that is sparse in the epistatic domain as simply being a “sparse fitness function”. A number of experimentally-determined fitness functions have been observed to be sparse in the epistatic domain [32, 17, 7]. Crucially, the sparsity of a fitness function in the epistatic domain determines how many measurements are required to estimate the fitness function using Compressed Sensing techniques that minimize a MSE loss function [7]. Herein we consider the effect that global epistasis has on a sparse fitness function. In particular, we argue that global epistasis results in observed fitness functions that are dense in the epistatic domain, and thus require a large amount of data to accurately estimate by minimizing a MSE loss function. However, in these cases, there may be a sparse ranking function that can be efficiently extracted by minimizing a contrastive loss function.
@@ 2.2 Global epistasis models
Global epistasis models of the form of Eq. 2 have proved effective at capturing the variation observed in empirical fitness data [26, 37, 31, 33, 34, 39], suggesting that global epistasis is a common feature of natural fitness functions. Further, it has been shown that global epistasis results from first-principles physical considerations that are common in many biological systems. In particular, Husain and Murugan [22] show that global epistasis arises when the physical dynamics of a system is dominated by slow, collective modes of motion, which is often the case for protein dynamics. Aside from intrinsic/endogenous sources, the process of measuring fitness can also introduce nonlinear effects that are dependent on the experiment and not on the underlying fitness function. For example, fitness data is often left-censored, as many sequence have fitness that falls below the detection threshold of an assay. Finally, global diminishing-returns epistatic patterns have been observed widely in both single and multi-gene settings where the interactions are among genes rather than within a gene [26, 34, 4].
Together, these results indicate that global epistasis is an effect that can be expected in empirically-observed fitness functions. In what follows, we argue that global epistasis is detrimental to effective modeling of fitness functions using standard techniques. In particular, global epistasis manifests itself by producing observed data that is dense in the epistatic domain. In other words, when observed fitness data is produced through Eq. 2 then the epistatic representation of this fitness function (calculated through application of Eq. 1), is not sparse. Further we argue that this effect of global epistasis makes it to difficult to model such observed data by minimizing standard MSE loss functions with a fixed amount of data. Further, we argue that fitting fitness models aimed at extracting the latent fitness function from observed data is a more efficient use of observed data that results in improved predictive performance (in the ranking sense).
@@ 2.3 Contrastive losses
where $f_{{\bm{\theta}}}$ is a model with parameters ${\bm{\theta}}$, ${\bm{x}}_{i}$ are model inputs and $y_{i}$ are the corresponding labels of those inputs. This loss compares every pair of data points and encourages the model output $f_{{\bm{\theta}}}({\bm{x}}_{i})$ to be greater than $f_{{\bm{\theta}}}({\bm{x}}_{j})$ whenever $y_{i}>y_{j}$; in other words, it encourages the model outputs to be ranked according to their labels. A number of distinct but similar loss functions have been proposed in the learning-to-rank literature [13] and also for metric learning [19]. An example is the Margin ranking loss [20], which replaces the logistic function in the sum of Eq. 3 with a hinge function. In our experiments, we focus on the BT loss of Eq. 3 as we found it typically results in superior predictive performance
@@
A key feature of the contrastive loss in Eq. 3 is that it only uses information about the ranking of observed labels, rather than the numerical values of the labels. Thus, the loss is unchanged when the observed values are transformed by a monotonic nonlinearity. We will show that this feature allows this loss to recover a sparse latent fitness function from observed data that has been affected by global epistasis, and enables more data-efficient learning of fitness functions compared to a MSE loss.
@@ 3 Results
Our results are aimed at demonstrating three properties of contrastive losses. First, we show that given complete, noiseless fitness data (i.e. noiseless fitness values associated with every sequence in the sequence space) that has been affected by global epistasis, minimizing the BT loss enables a model to nearly exactly recover the sparse latent fitness function $f$. Next, we consider the case of incomplete data, where the aim is to predict the relative fitness of unobserved sequences. In this regime, we find through simulation that minimizing the BT loss enables models to achieve better predictive performance then minimizing the MSE loss when the observed data has been affected with global epistasis. We argue by way of a fitness-epistasis uncertainty principle that this is due to the fact that nonlinearities tend to produce fitness functions that do not admit a sparse representation in the epistatic domain, and thus require more data to learn with MSE loss. Finally, we demonstrate the practical significance of these insights by showing that minimizing the BT loss results in improved performance over MSE loss in nearly all tasks in the FLIP benchmark [14].
@@ 3.1 Recovery from complete data
We first consider the case of “complete” data, where fitness measurements are available for every sequence in the sequence space. The aim of our task in this case is to recover a sparse latent fitness function when the observed measurements have been transformed by an arbitrary monotonic nonlinearity. In particular, we sample a sparse fitness function $f$ from the NK model [24], a popular model of fitness functions that has been shown to recapitulate the sparsity properties of some empirical fitness functions [7]. The NK model has three parameters: $L$, the length of the sequences, $q$, the size of the alphabet for sequence elements, and $K$, the maximum order of (local) epistatic interactions in the fitness function. Roughly, the model randomly assigns $K-1$ interacting positions to each position in the sequence, resulting in a sparse set of interactions in the epistatic domain. The weights of each of the assigned interactions are then drawn from a independent unit normal distributions.
@@
We then transform the sampled fitness function $f$ with a monotonic nonlinearity $g$ to produce a set of complete observed data, $y_{i}=g(f({\bm{x}}_{i}))$ for all ${\bm{x}}_{i}\in\mathcal{S}$. The goal of the task is then to recover the function $f$ given all $({\bm{x}}_{i},y_{i})$ pairs. In order to do so, we model $f$ using a two layer fully connected neural network and fit the parameters of this model by performing stochastic gradient descent (SGD) on the BT loss, using the Spearman correlation between model predictions and the $y_{i}$ values to determine convergence of the optimization. We then compare the resulting model, $\hat{f}$, to the latent fitness function $f$ in both the fitness and epistatic domains, using the forward and inverse transformation of Eq. 1 to convert between the two domains.
@@
Fig. 1 shows the results of one of these tests. In this case, we used an exponential function to represent the global epistasis nonlinearity. The exponential function exaggerates the effects of global epistasis in the epistatic domain and thus better illustrates the usefulness of contrastive losses, although the nonlinearities in empirical fitness functions tend to have a more sigmoidal shape [31]. Fig. 1b shows that the global epistasis nonlinearity substantially alters the representations of the observed data $y$ in both the fitness and epistatic domains, as compared to the latent fitness function $f$. Nonetheless, Fig. 1c demonstrates that the model fitness function $\hat{f}$ created by minimizing the BT loss is able to nearly perfectly recover the sparse latent fitness function (where recovery is defined as being equivalent up to an affine transformation). This is a somewhat surprising result, as there are many fitness functions that correctly rank the fitness of sequences, and it is not immediately clear why minimizing the BT loss produces this particular sparse latent fitness function. However, this example makes clear that fitting a model by minimizing the BT loss can be an effective strategy for recovering a sparse latent fitness function from observed data that has been affected by global epistasis. Similar results from additional examples of this task using different nonlinearities and latent fitness functions are shown in Appendix B.1.
@@ 3.2 Fitness-epistasis uncertainty principle
Figure 2: (a) Demonstration of the fitness-epistasis uncertainty principle for a latent fitness function transformed by $g(f)=\exp(a\cdot f)$ for various settings of $a$. The dashed black line indicates the lower bound on the sum of the entropies of the fitness and epistatic representations of the fitness function (b) Test-set Spearman correlation for models trained with MSE and BT losses on incomplete fitness data transformed by various nonlinearities, compared to the entropy of the fitness function in the epistatic domain. Each point corresponds to a model trained on 256 randomly sampled training points from an $L=10,K=2$ latent fitness function which was then transformed by a nonlinearity. (c) Convergence of models fit with BT and MSE losses to observed data generated by transforming an $L=10,K=2$ latent fitness function by $g(f)=\exp(10\cdot f)$. Each point represents the mean test set correlation over 200 training set replicates.
@@
Fig. 1b provides an example where a monotonic nonlinearity applied to a sparse fitness increases the density of the the fitness function in the epistatic domain. In particular, we see that many “spurious” local epistatic interactions must appear in order to represent the nonlinearity (e.g. interactions of order 3, when we used an NK model with $K=2$). This effect can be observed for many different shapes of nonlinearities [36, 3]. We can understand this effect more generally using uncertainty principles, which roughly show that a function cannot be concentrated on a small number of values in two different representations. In particular, we consider the discrete entropic uncertainty principle proved by Dembo et al. [15]. When applied to the transformation in Eq. 1, this uncertainty principle states:
$H({\bm{f}})+H(\bm{\beta})\geq-2L\log m,$ (4)
where $H({\bm{x}})=-\sum_{i}\frac{x^{2}_{i}}{||x||^{2}}\log\frac{x^{2}_{i}}{||x||^{2}}$ is the entropy of the normalized squared magnitudes of a vector and $m=1/\sqrt{q}$ when $q=2$, $m=1/(q-\sqrt{q})$ when $q=3$ and $m=1-1/(q-\sqrt{q})$ otherwise. Low entropy indicates that a vector is concentrated on a small set of elements. Thus, the fitness-epistasis uncertainty principle of Eq. 4 shows that fitness functions cannot be concentrated in both the fitness and epistatic domains. A sparse fitness function (in the epistatic domain) must therefore be “spread out” (i.e. dense) in the fitness domain, and vice-versa.
The importance of this result for understanding global epistasis is that applying a nonlinearity to a dense vector will often have the effect of concentrating the vector on a smaller number of values. This can most easily be seen for convex nonlinearities like the exponential shown in Fig. 1a, but is also true of many other nonlinearities (see Theorem 1 in Appendix C for a sufficient condition for a nonlinearity to decrease entropy in the fitness domain and Appendix D for additional experiments demonstrating the uncertainty principle). If this concentration in the fitness domain is sufficiently extreme, then the epistatic representation of the fitness function, $\bm{\beta}$, must be dense in order to satisfy Eq. 4. Fig. 2a demonstrates the uncertainty principle by showing how the entropy in the fitness and epistatic domains decrease and increase, respectively, as more extreme nonlinearities are applied to a sparse latent fitness function.